The length of a rectangle is three times the width. If the length is increased by 4 feet and the width is decreased by 1 foot, the area will be increased by 24 square feet. What are the dimensions of the original rectangle?
Answer by stanbon(48558)

The length of a rectangle is three times the width. If the length is increased by 4 feet and the width is decreased by 1 foot, the area will be increased by 24 square feet. What are the dimensions of the original rectangle?
------------
Original Rectangle:
Let width by "x"; length = 3x
Area = x*3x = 3x^2
-----------------------
Rectangle after change:
Width = x-1 ; length = 3x+4
Area = (x-1)(3x+4) = 3x^2+x-4
---------------
EQUATION:
original area + 24 = changed area
3x^2+24 = 3x^2+x-4
24 = x-4
x = 28 feet (original width)
3x = 84 feet (original length)
Cheers,
Stan H.